Salt is often used to melt ice and snow on roads in the winter and the temperature right now is -4.500ºC. How much NaCl (complete dissociation) would be necessary to melt 1000kg of ice by bringing down the freezing temperature to - 4.500ºC?


1
Expert's answer
2022-03-13T15:19:34-0400

(NaCl) = 58.5 g/mol;


Kf(H2O) = 1.86 (C * kg/mol)


\DeltaΔ Tf = b(molality) * Kf;


b = n(solution)/m(solvent) = m(NaCl)/(M(NaCl) * m(H2O));


m(NaCl), kg = b *M(NaCl) * m(H2O)/1000 = \DeltaΔ Tf *M(NaCl) * m(H2O) / (Kf * 1000)= 4.500 * 58.5 * 1000 / (1.86 *1000) = 141.5 kg.


Answer: 141.5 kg.